# RClicks: Realistic Click Simulation for Benchmarking Interactive Segmentation

### v1

@@ 7 Conclusion
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Introduction 1 and Conclusion 7.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix D.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix D.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
@@
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
@@
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We include it in supplemental material.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix A.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 4, Appendix C.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendices AC.
@@
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix D.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See the supplementary
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix B.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix B for an example of textual instruction.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Appendix B.
RClicks: Realistic Click Simulation
for Benchmarking Interactive Segmentation
Supplementary Materials
Appendix A Clickability Model
A.1 Training and model details
@@
Architecture details. As a predictor, we adapted the state-of-the-art SegNeXt-B segmentation network guo2022segnext with the MSCAN-B backbone. We input the original image into the network and concatenate the ground truth mask with the error mask, feeding the resulting tensor as an additional input to the network. This technique is inspired by Conv1S [ritm]. However, we used three convolutional layers with non-linear activation functions instead of one, as we needed to encode more complex features of the ground truth mask and error mask, rather than just clicks. Additionally, after a forward pass, we used min-max normalization to transform values to the interval from 0 to 1.
@@ A.2 Model generalizability
A.3 Clickability maps ablations
@@
A.4 Examples of generated clicks
@@
Appendix B Clicks Dataset
B.1 Collection procedure
@@
In determining the optimal display duration for our dataset collection, we conducted a thorough review of existing literature on annotation time requirements. A number of studies have been conducted on this subject  [delatolas2024learning] (Sec. 4),  [ravi2024sam] (Sec. E.1.2),  [Papadopoulos2014TrainingOC] (Sec. 2),  [papadopoulos2017training] (Sec. 3.4),  [bearman2016s] (Sec. 4.1). Researchers came to the conclusion that the required time plan formula is the following: 1 sec. for the annotator to visually locate an object and 1.5 sec. for adding each click, while total reported time to localize and click may vary between 1.87 sec. and 2.5 sec. Since, we did not have the goal of speeding up the labeling of RClicks, but to collect high-quality data, we increased times as follows: showing an image for 1.5 sec., different mask displaying modes for 2 sec. (since time is needed to localize and remember the target object) and Text Description mode for 2.5 sec. (since more time is needed to read and understand the text).
@@
B.2 Comparison PC and mobile clicks
@@
Appendix C Additional Benchmark Results
C.1 Evaluation setup
@@
C.2 Additional evaluation results
@@
Method Model Data GrabCut [grabcut] Berkeley [berkeley] DAVIS [davis] COCO-MVal [coco] TETRIS [moskalenko2024tetris]
@@
GPCIS [zhou2023interactivegpcis] RN50 C+L 2.04±0.49 1.80 13.58 61.59 20.38 3.06±0.83 2.61 22.13 57.13 24.53 6.44±0.85 5.94 16.88 53.65 19.31 4.74±1.31 4.05 26.43 79.00 30.02 3.87±0.79 3.39 19.55 56.43 22.90
CDNet [chen2021conditional] RN34 C+L 1.72±0.38 1.52 12.89 52.57 20.38 2.33±0.60 2.06 10.02 49.17 19.03 5.95±0.73 5.56 14.95 45.88 18.56 4.13±0.85 3.88 15.15 49.79 22.56 3.10±0.53 2.83 14.16 44.06 17.25
@@
CFR-ICL [sun2023cfricl] ViT-H C+L 1.49±0.32 1.58 0.32 24.23 11.26 1.57±0.27 1.46 6.11 23.58 7.69 4.53±0.46 4.24 9.32 18.47 9.21 2.70±0.63 2.70 9.58 24.13 12.83 2.12±0.34 2.01 8.76 14.33 8.68
MobileSAM [zhang2023faster] ViT-Tiny SA-1B 1.81±0.32 1.70 9.91 11.50 8.92 2.80±0.58 2.73 11.94 18.04 11.88 5.96±0.56 5.92 8.63 15.39 9.27 5.25±0.78 5.24 9.79 19.78 9.43 3.42±0.48 3.38 7.47 12.69 8.19
SAM [kirillov2023segment] ViT-B SA-1B 1.91±0.38 1.88 6.67 9.42 17.86 2.21±0.44 2.15 10.50 11.35 9.32 5.30±0.53 5.15 8.26 11.27 5.17 4.91±0.79 4.86 9.88 15.73 9.69 3.04±0.51 2.97 11.17 10.06 8.72
@@
Method Model Data GrabCut [grabcut] Berkeley [berkeley] DAVIS [davis] COCO-MVal [coco] TETRIS [moskalenko2024tetris]
@@
GPCIS [zhou2023interactivegpcis] RN50 C+L 0.0±0.00 0 0.0 0 0.0 1.5±0.81 1 0.5 1 0.6 61.0±3.82 59 2.0 11 4.8 46.2±15.89 48 -1.8 48 15.2 148.7±27.31 126 22.7 93 28.6
CDNet [chen2021conditional] RN34 C+L 0.0±0.00 0 0.0 0 0.0 0.9±0.94 0 0.9 3 1.4 61.8±5.10 56 5.8 16 6.0 58.0±5.85 58 0.0 15 6.4 144.6±14.85 126 18.6 48 21.6
@@
CFR-ICL [sun2023cfricl] ViT-H C+L 0.0±0.00 0 0.0 0 0.0 0.2±0.40 0 0.2 1 0.4 41.1±2.21 35 6.1 6 3.0 15.1±6.62 26 -10.9 17 7.0 62.4±4.29 58 4.4 15 5.6
MobileSAM [zhang2023faster] ViT-Tiny SA-1B 0.1±0.30 0 0.1 0 0.2 2.9±0.70 3 -0.1 0 0.2 65.3±1.00 67 -1.7 1 1.0 92.0±5.12 101 -9.0 15 1.2 181.4±5.82 178 3.4 13 1.2
SAM [kirillov2023segment] ViT-B SA-1B 0.0±0.00 0 0.0 0 0.0 0.9±0.30 1 -0.1 1 0.2 53.0±1.48 51 2.0 4 0.8 80.1±2.51 88 -7.9 0 -1.4 111.4±1.91 113 -1.6 2 -0.4
@@
Method Model Data GrabCut [grabcut] Berkeley [berkeley] DAVIS [davis] COCO-MVal [coco] TETRIS [moskalenko2024tetris]
@@
GPCIS [zhou2023interactivegpcis] RN50 C+L 97.16±0.83 97.29 -0.13 -1.71 -0.56 94.62±0.92 95.17 -0.55 -1.73 -0.61 89.30±1.27 90.39 -1.09 -2.86 -1.08 90.89±2.91 92.62 -1.73 -6.48 -2.30 92.73±1.83 93.98 -1.25 -3.69 -1.29
CDNet [chen2021conditional] RN34 C+L 96.58±1.59 97.14 -0.56 -3.80 -1.07 93.87±1.80 94.78 -0.91 -3.70 -1.39 88.05±2.05 89.71 -1.65 -5.18 -1.92 90.74±2.17 91.81 -1.08 -3.57 -1.43 92.98±1.97 94.19 -1.22 -3.97 -1.54
@@
CFR-ICL [sun2023cfricl] ViT-H C+L 98.66±0.57 98.35 0.31 -0.80 -0.38 96.57±0.60 96.70 -0.13 -0.93 -0.31 92.09±1.09 92.71 -0.62 -1.61 -0.66 94.54±1.26 94.84 -0.29 -1.71 -0.60 96.25±0.84 96.60 -0.36 -0.52 -0.19
MobileSAM [zhang2023faster] ViT-Tiny SA-1B 96.90±0.98 96.92 -0.02 -0.17 -0.12 93.39±1.36 93.56 -0.17 -0.11 -0.00 88.10±1.56 88.79 -0.69 -1.69 -0.64 88.64±2.58 89.26 -0.61 -0.79 -0.19 92.61±1.38 92.84 -0.23 -0.49 -0.14
SAM [kirillov2023segment] ViT-B SA-1B 96.97±0.93 96.99 -0.02 0.29 -0.19 94.52±0.86 94.57 -0.05 -0.02 -0.18 89.46±1.30 89.76 -0.30 -0.53 -0.14 90.70±1.53 91.03 -0.32 -0.23 -0.04 93.92±1.18 94.18 -0.26 0.09 0.04
@@
Method Model Data GrabCut [grabcut] Berkeley [berkeley] DAVIS [davis] COCO-MVal [coco] TETRIS [moskalenko2024tetris]
@@
GPCIS [zhou2023interactivegpcis] RN50 C+L 83.41±8.67 83.49±7.68 84.46 11.63 0.89 78.79±8.30 79.15±8.04 79.61 12.24 0.99 72.53±9.71 73.66±7.39 76.03 16.30 1.62 64.04±9.06 64.43±7.95 66.30 18.21 3.81 71.99±10.13 71.75±9.73 74.40 17.24 2.36
CDNet [chen2021conditional] RN34 C+L 87.85±10.22 88.27±8.19 90.80 12.50 1.66 81.92±10.42 82.47±9.45 84.64 14.69 1.96 71.97±11.50 73.41±8.24 76.70 19.72 2.53 76.05±10.28 76.80±8.00 79.22 16.32 2.65 79.56±11.38 79.61±10.84 82.47 17.32 2.31
@@
CFR-ICL [sun2023cfricl] ViT-H C+L 88.95±4.72 89.33±4.86 87.61 7.29 0.59 85.05±5.88 85.75±4.59 84.57 9.37 0.64 75.48±8.23 76.43±6.71 76.84 17.09 1.31 81.18±6.55 81.47±5.28 81.64 12.34 1.53 85.92±7.02 85.86±6.68 85.64 11.47 0.98
MobileSAM [zhang2023faster] ViT-Tiny SA-1B 76.61±14.67 77.02±13.49 79.14 24.45 1.04 75.30±15.97 76.03±14.91 76.83 27.28 1.55 58.17±16.01 59.48±14.55 61.87 40.95 2.08 66.65±15.01 66.85±13.09 65.53 32.55 3.30 66.81±18.77 67.12±18.60 65.72 36.21 1.78
SAM [kirillov2023segment] ViT-B SA-1B 64.65±21.07 63.52±20.47 62.43 42.61 0.97 66.02±20.18 66.81±18.58 68.47 41.66 0.92 45.27±17.48 46.55±16.41 48.55 57.06 1.59 60.90±18.20 61.20±15.88 58.29 44.16 1.84 55.05±21.59 54.99±21.25 55.88 51.19 1.38
@@
C.3 Finding hard instances for the first round using real-user clicks.
@@
Figures 15 and 16 illustrate the hard (high NSR) and simple (low NSR) cases for RITM HRNet32-IT (C+L) ritm, SimpleClick ViT-H (C+L) liu2022simpleclick, SAM ViT-H (SA1-B) kirillov2023segment, SAM-HQ ViT-H (SA1-B) ke2023segment. For each method, the uncertainty of prediction is depicted by the mean mask averaged over the method’s predictions on the first-round clicks. In the averaged mask, gray pixels correspond to variations of the method predictions (hard cases), black and white pixels illustrate low variation in, respectively, background and foreground predictions (simple cases).
Figure 15: Samples with high User-IoU@1-NSR. From left to right – the image with the target instance and real-user clicks of the first round; masks, averaged over the clicks, obtained by the methods: RITM HRNet32-IT (C+L) ritm, SimpleClick ViT-H (C+L) liu2022simpleclick, SAM ViT-H (SA1-B) kirillov2023segment, SAM-HQ ViT-H (SA1-B) ke2023segment.
Figure 16: Samples with low User-IoU@1-NSR. From left to right – the image with the target instance and real-user clicks of the first round; masks, averaged over the clicks, obtained by the methods: RITM HRNet32-IT (C+L) ritm, SimpleClick ViT-H (C+L) liu2022simpleclick, SAM ViT-H (SA1-B) kirillov2023segment, SAM-HQ ViT-H (SA1-B) ke2023segment.
Appendix D Benchmark Discussion
  1. 1.
  2. 2.
  3. 3.
  4. 4.

### v2

@@ Appendix A Benchmark Discussion
Appendix B Clickability Model
B.1 Training and model details
@@
Architecture details. As a predictor, we adapted the state-of-the-art SegNeXt-B segmentation network Guo et al. [2022] with the MSCAN-B backbone. We input the original image into the network and concatenate the ground truth mask with the error mask, feeding the resulting tensor as an additional input to the network. This technique is inspired by Conv1S [Sofiiuk et al., 2022]. However, we used three convolutional layers with non-linear activation functions instead of one, as we needed to encode more complex features of the ground truth mask and error mask, rather than just clicks. Additionally, after a forward pass, we used min-max normalization to transform values to the interval from 0 to 1.
@@ B.2 Model generalizability
B.3 Clickability maps ablations
@@
B.4 Examples of generated clicks
@@
Appendix C Clicks Dataset
C.1 Collection procedure
@@
In determining the optimal display duration for our dataset collection, we conducted a thorough review of existing literature on annotation time requirements. A number of studies have been conducted on this subject  [Delatolas et al., 2024] (Sec. 4),  [Ravi et al., 2024] (Sec. E.1.2),  [Papadopoulos et al., 2014] (Sec. 2),  [Papadopoulos et al., 2017] (Sec. 3.4),  [Bearman et al., 2016] (Sec. 4.1). Researchers came to the conclusion that the required time plan formula is the following: 1 sec. for the annotator to visually locate an object and 1.5 sec. for adding each click, while total reported time to localize and click may vary between 1.87 sec. and 2.5 sec. Since, we did not have the goal of speeding up the labeling of RClicks, but to collect high-quality data, we increased times as follows: showing an image for 1.5 sec., different mask displaying modes for 2 sec. (since time is needed to localize and remember the target object) and Text Description mode for 2.5 sec. (since more time is needed to read and understand the text).
@@
C.2 Comparison of PC and mobile clicks
@@
Appendix D Additional Benchmark Results
D.1 Evaluation setup
@@
D.2 Additional evaluation results
@@
Method Model Data GrabCut [Rother et al., 2004] Berkeley [Martin et al., 2001] DAVIS [Perazzi et al., 2016] COCO-MVal [Lin et al., 2014a] TETRIS [Moskalenko et al., 2024]
@@
GPCIS [Zhou et al., 2023b] RN50 C+L 2.04±0.49 1.80 13.58 61.59 20.38 3.06±0.83 2.61 22.13 57.13 24.53 6.44±0.85 5.94 16.88 53.65 19.31 4.74±1.31 4.05 26.43 79.00 30.02 3.87±0.79 3.39 19.55 56.43 22.90
CDNet [Chen et al., 2021] RN34 C+L 1.72±0.38 1.52 12.89 52.57 20.38 2.33±0.60 2.06 10.02 49.17 19.03 5.95±0.73 5.56 14.95 45.88 18.56 4.13±0.85 3.88 15.15 49.79 22.56 3.10±0.53 2.83 14.16 44.06 17.25
@@
CFR-ICL [Sun et al., 2023] ViT-H C+L 1.49±0.32 1.58 0.32 24.23 11.26 1.57±0.27 1.46 6.11 23.58 7.69 4.53±0.46 4.24 9.32 18.47 9.21 2.70±0.63 2.70 9.58 24.13 12.83 2.12±0.34 2.01 8.76 14.33 8.68
MobileSAM [Zhang et al., 2023] ViT-Tiny SA-1B 1.81±0.32 1.70 9.91 11.50 8.92 2.80±0.58 2.73 11.94 18.04 11.88 5.96±0.56 5.92 8.63 15.39 9.27 5.25±0.78 5.24 9.79 19.78 9.43 3.42±0.48 3.38 7.47 12.69 8.19
SAM [Kirillov et al., 2023] ViT-B SA-1B 1.91±0.38 1.88 6.67 9.42 17.86 2.21±0.44 2.15 10.50 11.35 9.32 5.30±0.53 5.15 8.26 11.27 5.17 4.91±0.79 4.86 9.88 15.73 9.69 3.04±0.51 2.97 11.17 10.06 8.72
@@
Method Model Data GrabCut [Rother et al., 2004] Berkeley [Martin et al., 2001] DAVIS [Perazzi et al., 2016] COCO-MVal [Lin et al., 2014a] TETRIS [Moskalenko et al., 2024]
@@
GPCIS [Zhou et al., 2023b] RN50 C+L 0.0±0.00 0 0.0 0 0.0 1.5±0.81 1 0.5 1 0.6 61.0±3.82 59 2.0 11 4.8 46.2±15.89 48 -1.8 48 15.2 148.7±27.31 126 22.7 93 28.6
CDNet [Chen et al., 2021] RN34 C+L 0.0±0.00 0 0.0 0 0.0 0.9±0.94 0 0.9 3 1.4 61.8±5.10 56 5.8 16 6.0 58.0±5.85 58 0.0 15 6.4 144.6±14.85 126 18.6 48 21.6
@@
CFR-ICL [Sun et al., 2023] ViT-H C+L 0.0±0.00 0 0.0 0 0.0 0.2±0.40 0 0.2 1 0.4 41.1±2.21 35 6.1 6 3.0 15.1±6.62 26 -10.9 17 7.0 62.4±4.29 58 4.4 15 5.6
MobileSAM [Zhang et al., 2023] ViT-Tiny SA-1B 0.1±0.30 0 0.1 0 0.2 2.9±0.70 3 -0.1 0 0.2 65.3±1.00 67 -1.7 1 1.0 92.0±5.12 101 -9.0 15 1.2 181.4±5.82 178 3.4 13 1.2
SAM [Kirillov et al., 2023] ViT-B SA-1B 0.0±0.00 0 0.0 0 0.0 0.9±0.30 1 -0.1 1 0.2 53.0±1.48 51 2.0 4 0.8 80.1±2.51 88 -7.9 0 -1.4 111.4±1.91 113 -1.6 2 -0.4
@@
Method Model Data GrabCut [Rother et al., 2004] Berkeley [Martin et al., 2001] DAVIS [Perazzi et al., 2016] COCO-MVal [Lin et al., 2014a] TETRIS [Moskalenko et al., 2024]
@@
GPCIS [Zhou et al., 2023b] RN50 C+L 97.16±0.83 97.29 -0.13 -1.71 -0.56 94.62±0.92 95.17 -0.55 -1.73 -0.61 89.30±1.27 90.39 -1.09 -2.86 -1.08 90.89±2.91 92.62 -1.73 -6.48 -2.30 92.73±1.83 93.98 -1.25 -3.69 -1.29
CDNet [Chen et al., 2021] RN34 C+L 96.58±1.59 97.14 -0.56 -3.80 -1.07 93.87±1.80 94.78 -0.91 -3.70 -1.39 88.05±2.05 89.71 -1.65 -5.18 -1.92 90.74±2.17 91.81 -1.08 -3.57 -1.43 92.98±1.97 94.19 -1.22 -3.97 -1.54
@@
CFR-ICL [Sun et al., 2023] ViT-H C+L 98.66±0.57 98.35 0.31 -0.80 -0.38 96.57±0.60 96.70 -0.13 -0.93 -0.31 92.09±1.09 92.71 -0.62 -1.61 -0.66 94.54±1.26 94.84 -0.29 -1.71 -0.60 96.25±0.84 96.60 -0.36 -0.52 -0.19
MobileSAM [Zhang et al., 2023] ViT-Tiny SA-1B 96.90±0.98 96.92 -0.02 -0.17 -0.12 93.39±1.36 93.56 -0.17 -0.11 -0.00 88.10±1.56 88.79 -0.69 -1.69 -0.64 88.64±2.58 89.26 -0.61 -0.79 -0.19 92.61±1.38 92.84 -0.23 -0.49 -0.14
SAM [Kirillov et al., 2023] ViT-B SA-1B 96.97±0.93 96.99 -0.02 0.29 -0.19 94.52±0.86 94.57 -0.05 -0.02 -0.18 89.46±1.30 89.76 -0.30 -0.53 -0.14 90.70±1.53 91.03 -0.32 -0.23 -0.04 93.92±1.18 94.18 -0.26 0.09 0.04
@@
Method Model Data GrabCut [Rother et al., 2004] Berkeley [Martin et al., 2001] DAVIS [Perazzi et al., 2016] COCO-MVal [Lin et al., 2014a] TETRIS [Moskalenko et al., 2024]
@@
GPCIS [Zhou et al., 2023b] RN50 C+L 83.41±8.67 83.49±7.68 84.46 11.63 0.89 78.79±8.30 79.15±8.04 79.61 12.24 0.99 72.53±9.71 73.66±7.39 76.03 16.30 1.62 64.04±9.06 64.43±7.95 66.30 18.21 3.81 71.99±10.13 71.75±9.73 74.40 17.24 2.36
CDNet [Chen et al., 2021] RN34 C+L 87.85±10.22 88.27±8.19 90.80 12.50 1.66 81.92±10.42 82.47±9.45 84.64 14.69 1.96 71.97±11.50 73.41±8.24 76.70 19.72 2.53 76.05±10.28 76.80±8.00 79.22 16.32 2.65 79.56±11.38 79.61±10.84 82.47 17.32 2.31
@@
CFR-ICL [Sun et al., 2023] ViT-H C+L 88.95±4.72 89.33±4.86 87.61 7.29 0.59 85.05±5.88 85.75±4.59 84.57 9.37 0.64 75.48±8.23 76.43±6.71 76.84 17.09 1.31 81.18±6.55 81.47±5.28 81.64 12.34 1.53 85.92±7.02 85.86±6.68 85.64 11.47 0.98
MobileSAM [Zhang et al., 2023] ViT-Tiny SA-1B 76.61±14.67 77.02±13.49 79.14 24.45 1.04 75.30±15.97 76.03±14.91 76.83 27.28 1.55 58.17±16.01 59.48±14.55 61.87 40.95 2.08 66.65±15.01 66.85±13.09 65.53 32.55 3.30 66.81±18.77 67.12±18.60 65.72 36.21 1.78
SAM [Kirillov et al., 2023] ViT-B SA-1B 64.65±21.07 63.52±20.47 62.43 42.61 0.97 66.02±20.18 66.81±18.58 68.47 41.66 0.92 45.27±17.48 46.55±16.41 48.55 57.06 1.59 60.90±18.20 61.20±15.88 58.29 44.16 1.84 55.05±21.59 54.99±21.25 55.88 51.19 1.38
@@
D.3 Finding hard instances for the first round using real-user clicks
@@
Figures 15 and 16 illustrate the hard (high NSR) and simple (low NSR) cases for RITM HRNet32-IT (C+L) Sofiiuk et al. [2022], SimpleClick ViT-H (C+L) Liu et al. [2022], SAM ViT-H (SA1-B) Kirillov et al. [2023], SAM-HQ ViT-H (SA1-B) Ke et al. [2023]. For each method, the uncertainty of prediction is depicted by the mean mask averaged over the method’s predictions on the first-round clicks. In the averaged mask, gray pixels correspond to variations of the method predictions (hard cases), black and white pixels illustrate low variation in, respectively, background and foreground predictions (simple cases).
Figure 15: Samples with high User-IoU@1-NSR. From left to right – the image with the target instance and real-user clicks of the first round; masks, averaged over the clicks, obtained by the methods: RITM HRNet32-IT (C+L) Sofiiuk et al. [2022], SimpleClick ViT-H (C+L) Liu et al. [2022], SAM ViT-H (SA1-B) Kirillov et al. [2023], SAM-HQ ViT-H (SA1-B) Ke et al. [2023].
Figure 16: Samples with low User-IoU@1-NSR. From left to right – the image with the target instance and real-user clicks of the first round; masks, averaged over the clicks, obtained by the methods: RITM HRNet32-IT (C+L) Sofiiuk et al. [2022], SimpleClick ViT-H (C+L) Liu et al. [2022], SAM ViT-H (SA1-B) Kirillov et al. [2023], SAM-HQ ViT-H (SA1-B) Ke et al. [2023].